OpenStudy (anonymous):

If w = <-8, 15>, find ||w||.

OpenStudy (zzr0ck3r):

\(||\langle a, b \rangle ||=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

@zzr0ck3r is a=-8, b=15?

OpenStudy (anonymous):

I got sqrt161, but tat's not my option.

OpenStudy (zzr0ck3r):

\(\sqrt{15^2+8^2} = \sqrt{289}\)